How much positive and negative charge is there in a cup of water

Grade:12

3 Answers

Sujit Kumar
6 years ago
Let us consider 1 cup = 250 mL = 250g
We know Density of Water = 1 \ \frac{g}{cc} (or) 1 \ \frac{g}{mL^3}
where g=grams, cc= centimeter cube, mL= mili litre
 
We know the formula of water is {{H}__2}O
Therefore in 18g of {{H}__2}O there are 8 protons and 8 electrons.
=> The number of protons and electrons each in 250 g of water are \frac{8\ast250}{18}\approx 111
=> There are 111 protons and 111 electrons in 250 mL of water
We know charge of 1 electron = -1.6\ast10^{-19} \ C and
charge of 1 proton =  +1.6\ast10^{-19} \ C
 
Therefore 
positive charge in a cup of water = 111\ast(+1.6\ast10^{-19})\approx+118\ast10^{-19}=+1.18*10^{-17} C
and
positive charge in cup of water = 111\ast(-1.6\ast10^{-19})\approx-118\ast10^{-19}=-1.18*10^{-17} C
Divya
6 years ago
Mass of water 250gMolar mass 18gNo of molecules in 18g 6.02*10^23No of molecules in one cup of water=250/18×6.02*10^23Each molecule of water contains 2Hydrogen and 1Oxygen i.e,10 electron and 10 protonHence Total positive charge in one cup=250/18×6.02×10^23×10×1.6×10^-19C=1.34×10^7C.Therefore,Total charge=2×1.34*10^7C =2.68C
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Assume that one cup holds 250 ml of water.
250 ml = 250 gram of water since its density is 1 in cgs.
number of moles of water in 250 gram = 250 / molar weight of water
n = 250/18 = 13.88 moles
1 mol has 6.022 x 1023 number of water molecules
So 250 g water will have 13.88 x 6.022 x 1023 number of H2O molecules
= 8.36 x 1024 number of molecules
One molecule of water has 2 hydrogen atoms (2 protons) and one oxygen atom (8 protons).
Total protons = 10 = total electrons (since it is neutral)
So 250 g water will have 10 x 8.36 x 1024 = 8.36 x 1025 protons.
charge of proton = 1.6 x 10-19 C
So total positive charge of 250 g water = 1.6 x 10-19 x 8.36 x 1025 = 13.37 x 106 C
which is equal to 13.3 MC (mega coulomb)
So one cup of water has 13.3 mega coulomb positive and negative charge, so that the overall charge is zero...

Thanks and Regards